OpenStudy (mrs.irwin):

Katelyn invested $2325 in a savings account which earns 3.45% interest compounded monthly. What will the account be worth in 7 years? $2,959.07 $25,144.92 $2,372.20 $40,159.15

OpenStudy (amistre64):

do you have your compounding formula ready?

OpenStudy (mrs.irwin):

@amistre64 that one is A = P (1 + r/n) ^ nt right?

OpenStudy (amistre64):

yes, thats the perfect one to use

OpenStudy (mrs.irwin):

@amistre64 so what do I plug in? I am not good in math

OpenStudy (amistre64):

even if we mess up and forgo the n parts, we are only 10 buck off ....

OpenStudy (amistre64):

plugging in isnt math, its more like reading comprehension

OpenStudy (mrs.irwin):

lol im not good at the either

OpenStudy (amistre64):

P is the amount you start with r is the rate in decimal form n is the number of compounding periods in a year t is the number of years

OpenStudy (amistre64):

forget n at the moment ... its really not worth it in this setup

OpenStudy (mrs.irwin):

okay so p= $2325 r= 3.45 and t= 7

OpenStudy (amistre64):

yes, let r = .0345 unless your calculator does % stuff. i never could get that utton to do what it was spose to

OpenStudy (mrs.irwin):

Now it looks like A= 2325(1+.0345/n)^n*7

OpenStudy (amistre64):

let n=1 or n=12; your choice .... but n=1 isnt going to be that far off

OpenStudy (mrs.irwin):

A=2325(1.0345)^7 right?

OpenStudy (amistre64):

just for extra stuff to type ... compare the results of 1 year 2325(1.0345) to 2325(1+.0345/12) 1.0345 : 1+.0345/12 1.0345 : 1.002875 1 : .9695 they are practically the same, almost a 1:1 ratio

OpenStudy (amistre64):

yes

OpenStudy (mrs.irwin):

okay i got 2948.06 but that isn't an answer

OpenStudy (amistre64):

youre not looking for THAT answer, your looking for the one thats closest to it ....

OpenStudy (mrs.irwin):

A is the closest to it

OpenStudy (amistre64):

yes it is :) now if they had all been within a 10 dollar range, then we would have had to go more precise ... but as is the choices wernt really that picky to start with

OpenStudy (mrs.irwin):

Thank you so much!!

OpenStudy (amistre64):

good luck